Last week we vented and we cleaned the main optics of the arm cavities.
I measured the frequency of the cavity poles for both the arm cavities to see how they changed (see previous elog entry 2226). These the results:
fp_X = 1616 +/- 14 Hz
fp_Y = 1590 +/- 4 Hz